# Proximal Interacting Particle Langevin Algorithms

### v1

@@ Abstract
We introduce a class of algorithms, termed Proximal Interacting Particle Langevin Algorithms (PIPLA), for inference and learning in latent variable models whose joint probability density is non-differentiable. Leveraging proximal Markov chain Monte Carlo (MCMC) techniques and the recently introduced interacting particle Langevin algorithm (IPLA), we propose several variants within the novel proximal IPLA family, tailored to the problem of estimating parameters in a non-differentiable statistical model. We prove nonasymptotic bounds for the parameter estimates produced by multiple algorithms in the strongly log-concave setting and provide comprehensive numerical experiments on various models to demonstrate the effectiveness of the proposed methods. In particular, we demonstrate the utility of the proposed family of algorithms on a toy hierarchical example where our assumptions can be checked, as well as on the problems of sparse Bayesian logistic regression, sparse Bayesian neural network, and sparse matrix completion. Our theory and experiments together show that PIPLA family can be the de facto choice for parameter estimation problems in latent variable models for non-differentiable models.
@@ 1 Introduction
This paper is organised as follows. In Section 2, we present the technical background necessary to develop our methods. In Section 3, we introduce the proposed algorithms—we provide their theoretical analysis in Section 4. Finally, in Section 5, we show comprehensive numerical experiments and we conclude with Section 6.
@@ 1.1 Notation
$W_{2}(\mu,\nu)=\left(\inf_{\gamma\in\mathbf{T}(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{2}\mathrm{d}\gamma(x,y)\right)^{1/2}$
where $\mathbf{T}(\mu,\nu)=\{\gamma\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d}):\gamma(A\times\mathbb{R}^{d})=\mu(A),\ \gamma(\mathbb{R}^{d}\times A)=\nu(A)\ \forall A\in\mathcal{B}(\mathbb{R}^{d})\}$ denotes the set of all transport plans between $\mu$ and $\nu$. In the following, we metrise $\mathcal{P}_{2}(\mathbb{R}^{d})$ with $W_{2}$.
@@ 2.1 Langevin Dynamics
where $\xi_{n}=\mathbf{B}_{n\gamma}-\mathbf{B}_{(n-1)\gamma}$ is a $d$-dimensional standard Gaussian. In the sampling literature, the algorithm (3) is known as the unadjusted Langevin algorithm (ULA) [Durmus and Moulines, 2019]. When $U$ is $\mu$-strongly convex and $L$-smooth (i.e. $\nabla U(x)$ is $L$-Lipschitz), ULA has very favourable properties. In particular, it can be shown that ULA converges to its own (biased) limit $\pi^{\gamma}$ exponentially fast, and that the asymptotic bias is of order $\gamma^{1/2}$ [Durmus and Moulines, 2019]. This makes ULA a favourable alternative to the Metropolis-adjusted Langevin algorithm (MALA) - which is asymptotically exact - because ULA does not require the computation of acceptance ratios and is cheaper in high dimensions [Durmus and Moulines, 2019].
@@ 2.2 MMLE with Langevin Dynamics
$\bar{\theta}_{\star}\in\arg\max_{\theta\in\Theta}p_{\theta}(y)=\arg\max_{\theta\in\Theta}\int_{\mathbb{R}^{d_{x}}}p_{\theta}(x,y)\mathrm{d}x.$ (4)
@@
where $(\xi_{n}^{i,N})_{n\in\mathbb{N}}$ are i.i.d. standard Gaussian random variables for all $i=0,1,\dots,N$. Under Lipschitz continuity and strong-convexity of $U$, IPLA has been shown to enjoy favourable convergence properties [Akyildiz et al., 2023].
@@ 2.3 Proximal Markov chain Monte Carlo
$\mathrm{d}\bm{X}_{\lambda,t}=\nabla\log\pi_{\lambda}(\bm{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t},$ (7)
@@
$\mathrm{d}\bm{X}_{\lambda,t}=\lambda^{-1}(\operatorname{prox}_{U}^{\lambda}(\bm{X}_{\lambda,t})-\bm{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}.$
@@ 2.3.1 Proximal Langevin methods
$X_{n+1}=\Big{(}1-\frac{\gamma}{\lambda}\Big{)}X_{n}+\frac{\gamma}{\lambda}\operatorname{prox}_{U}^{\lambda}(X_{n})+\sqrt{2\gamma}\,\xi_{n+1},$ (8)
where $\xi_{n}=\mathbf{B}_{n\gamma}-\mathbf{B}_{(n-1)\gamma}$, for $n\geq 0$, are $d$-dimensional standard Gaussians.
@@
for $n\geq 0$, where $\xi_{n}$ are i.i.d. standard Gaussians. The computational performance of the algorithm depends strongly on the capacity to evaluate efficiently $\operatorname{prox}_{U}^{\gamma}$.
@@
$\mathrm{d}\bm{X}_{\lambda,t}=-(\nabla g_{1}(\bm{X}_{\lambda,t})+\nabla g_{2}^{\lambda}(\bm{X}_{\lambda,t}))\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t},$
@@
$\mathrm{d}\bm{X}_{\lambda,t}=-\nabla g_{1}(\bm{X}_{\lambda,t})+\lambda^{-1}(\operatorname{prox}_{g_{2}}^{\lambda}(\bm{X}_{\lambda,t})-\bm{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}.$
@@ 2.3.2 Proximal gradient MCMC methods
where $\xi_{n}=\mathbf{B}_{n\gamma}-\mathbf{B}_{(n-1)\gamma}$ are $d$-dimensional standard Gaussians and $\gamma$ is the time discretisation step. Combining all the steps we obtain the proximal gradient Langevin algorithm (PGLA)
@@ 3 Proximal Interacting Particle Methods for MMLE
$\operatorname{prox}_{U}^{\lambda}(\theta,x)=(\operatorname{prox}_{U}^{\lambda}(\theta,x)_{\theta},\operatorname{prox}_{U}^{\lambda}(\theta,x)_{x})=\operatorname*{arg\,min}_{z_{0}\in\mathbb{R}^{d_{\theta}},z\in\mathbb{R}^{d_{x}}}\;\left\{U(z_{0},z)+\|(z_{0},z)-(\theta,x)\|^{2}/(2\lambda)\right\}.$
@@ 4.1 Assumptions
$\langle v-v^{\prime},\nabla U^{\lambda}(v)-\nabla U^{\lambda}(v^{\prime})\rangle=\langle v-v^{\prime},\nabla g_{1}(v)-\nabla g_{1}(v^{\prime})\rangle+\langle v-v^{\prime},\nabla g_{2}^{\lambda}(v)-\nabla g_{2}^{\lambda}(v^{\prime})\rangle\geq\mu\|v-v^{\prime}\|^{2}.$
@@ 4.2 The proof strategy
$\pi^{N}_{\lambda,\Theta}(\theta)\propto\int_{\mathbb{R}^{d_{x}}}...\int_{\mathbb{R}^{d_{x}}}e^{-\sum^{N}_{i=1}U^{\lambda}(\theta,x_{i})}\mathrm{d}x_{1}\mathrm{d}x_{2}\dots\mathrm{d}x_{N}=\biggr{(}\int_{\mathbb{R}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x\biggr{)}^{N}.$
@@
The first term is a concentration result, characterising the concentration of the $\theta$-marginal of the target measure of the SDE (9)–(10) around the maximiser of $p_{\theta}(y)$. Handling this term is not trivial since the maximisers of $p_{\theta}(y)$ and of $p_{\theta}^{\lambda}(y):=\int_{\mathbb{R}^{d_{x}}}p_{\theta}^{\lambda}(x,y)\mathrm{d}x$ are not necessarily the same as we clarify in the next section. The second term is a convergence term, which characterises the convergence of the solution of the SDE to its target measure. The last term is a discretisation term, which characterises the error introduced by discretising the SDE. By separately bounding each term, we provide nonasymptotic results for the convergence of the proximal interacting particle algorithms. Proofs of the following results are provided in Appendices A and B.
@@ 4.3 Nonasymptotic analysis of MYIPLA
$\mathbb{E}[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq$ $\frac{\lambda}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}+\sqrt{\frac{2d_{\theta}}{N\mu}}+e^{-\mu n\gamma}\bigg{(}\mathbb{E}[\|Z_{0}^{N}-z_{\star}\|^{2}]^{1/2}+\Big{(}\frac{d_{x}N+d_{\theta}}{N\mu}\Big{)}^{1/2}\bigg{)}$
@@
$\|\bar{\theta}_{\star}-\bar{\theta}_{\star,\lambda}\|\leq\frac{\lambda}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}+\mathcal{O}(\lambda^{2}),$
with $A,B$ given in A5. This term quantifies the distance between maximisers of $p_{\theta}(y)$ and $p^{\lambda}_{\theta}(y)$. Next, consider the stationary measure of the SDE (9)–(10) denoted by $\pi_{\lambda,\star}^{N}(\theta,x_{1},\ldots,x_{N})$. We are interested in the $\theta$-marginal of this measure, denoted by $\pi_{\lambda,\Theta}^{N}$. Using Proposition 3 of Akyildiz et al. [2023], we have that
$W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\star}})\leq\sqrt{\frac{2d_{\theta}}{N\mu}},$
@@ 4.4 Nonasymptotic analysis of PIPGLA
To derive a convergence rate for PIPGLA we make the following additional assumption
@@
For any given $\lambda>0$ we assume that $\mathbb{E}[\|\nabla g_{2}^{\lambda}(\theta^{N}_{n+1/2},X_{n+1}^{i,N})\|^{2}]\leq C$ for all $n\in\mathbb{N}$, $i=1,\dots,N$.
This assumption is similar to Salim et al. [2019, Assumption 5] and guarantees that the second moment of $\nabla g_{2}^{\lambda}$ does not explode.
@@
$\mathbb{E}\left[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}\right]^{1/2}\leq$ $\frac{\lambda}{\mu}\left(\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\right)+\sqrt{\frac{2d_{\theta}}{N\mu}}+\frac{\lambda^{n/2}(1-\gamma\mu)^{n/2}}{\gamma^{n/2}}W_{2}(\mathcal{L}({Z_{0}^{N}}),\pi_{\lambda}^{N})$ (20)
$+\left(\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda C)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}\right)^{1/2}+\mathcal{O}(\lambda^{2}),$
@@
Similarly to the previous result, we split the errors into same terms as given in (19). The concentration term in (19) is bounded as before
$W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\lambda,\Theta}^{N})\leq\frac{\lambda}{\mu}\left(\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\right)+\sqrt{\frac{2d_{\theta}}{N\mu}}+\mathcal{O}(\lambda^{2}).$
In the case of PIPGLA, we derive novel nonasymptotic bounds for the convergence and discretisation error (see Corollary A.6 in Appendix A.4). Intuitively, the third term in (20) controls the convergence to $\pi_{\lambda}^{N}$, while the fourth only acts on the error due to time discretisation.

### v2

@@ Abstract
We introduce a class of algorithms, termed Proximal Interacting Particle Langevin Algorithms (PIPLA), for inference and learning in latent variable models whose joint probability density is non-differentiable. Leveraging proximal Markov chain Monte Carlo (MCMC) techniques and the recently introduced interacting particle Langevin algorithm (IPLA), we propose several variants within the novel proximal IPLA family, tailored to the problem of estimating parameters in a non-differentiable statistical model. We prove nonasymptotic bounds for the parameter estimates produced by the different algorithms in the strongly log-concave setting and provide comprehensive numerical experiments on various models to demonstrate the effectiveness of the proposed methods. In particular, we demonstrate the utility of our family of algorithms on a toy hierarchical example where our assumptions can be checked, as well as for sparse Bayesian logistic regression, training of sparse Bayesian neural networks, and sparse matrix completion. Our theory and experiments together show that PIPLA family can be the de facto choice for parameter estimation problems in non-differentiable latent variable models.
@@ 1 Introduction
This paper is organised as follows. In Section 2, we present the technical background necessary to develop our methods. In Section 3, we introduce the proposed algorithms—we provide their theoretical analysis in Section 4. Finally, in Section 5, we show comprehensive numerical experiments and we conclude with Section 6. The code is available in the GitHub repository: https://github.com/paulaoak/proximal-ipla.
@@ 1.1 Notation
$W_{2}(\mu,\nu)=\left(\inf_{\gamma\in\mathbf{T}(\mu,\nu)}\int_{{}^{d}\times^{d}}\|x-y\|^{2}\mathrm{d}\gamma(x,y)\right)^{1/2}$
where $\mathbf{T}(\mu,\nu)=\{\gamma\in\mathcal{P}(^{d}\times^{d}):\gamma(A\times^{d})=\mu(A),\ \gamma(^{d}\times A)=\nu(A)\ \forall A\in\mathcal{B}(^{d})\}$ denotes the set of all transport plans between $\mu$ and $\nu$. In the following, we metrise $\mathcal{P}_{2}(^{d})$ with $W_{2}$.
@@ 2.1 Langevin Dynamics
where $(\xi_{n})_{n\in\mathbb{N}}$ is a sequence of i.i.d. $d$-dimensional standard Gaussian random variables. In the sampling literature, the algorithm (3) is known as the unadjusted Langevin algorithm (ULA) [Durmus and Moulines, 2019]. When $U$ is $\mu$-strongly convex and $L$-smooth (i.e. $\nabla U(x)$ is $L$-Lipschitz), ULA has very favourable properties. In particular, it can be shown that ULA converges to its own (biased) limit $\pi^{\gamma}$ exponentially fast, and that the asymptotic bias is of order $\gamma^{1/2}$ [Durmus and Moulines, 2019]. This makes ULA a favourable alternative to the Metropolis-adjusted Langevin algorithm (MALA) - which is asymptotically exact - because ULA does not require the computation of acceptance ratios and is cheaper in high dimensions [Durmus and Moulines, 2019].
@@ 2.2 MMLE with Langevin Dynamics
$\bar{\theta}_{\star}\in\arg\max_{\theta\in\Theta}p_{\theta}(y)=\arg\max_{\theta\in\Theta}\int_{{}^{d_{x}}}p_{\theta}(x,y)\mathrm{d}x.$ (4)
@@
where $(\xi_{n}^{i,N})_{n\in\mathbb{N}}$ are i.i.d. standard Gaussian random variables of the corresponding dimension for all $i=0,1,\dots,N$. Under Lipschitz continuity and strong-convexity of $U$, IPLA has been shown to enjoy favourable convergence properties [Akyildiz et al., 2023].
@@ 2.3 Proximal Markov chain Monte Carlo
$\mathrm{d}\mathbf{X}_{\lambda,t}=\nabla\log\pi_{\lambda}(\mathbf{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t},$ (7)
@@
$\mathrm{d}\mathbf{X}_{\lambda,t}=\lambda^{-1}(\operatorname{prox}_{U}^{\lambda}(\mathbf{X}_{\lambda,t})-\mathbf{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}.$
@@ 2.3.1 Proximal Langevin methods
$X_{n+1}=\Big{(}1-\frac{\gamma}{\lambda}\Big{)}X_{n}+\frac{\gamma}{\lambda}\operatorname{prox}_{U}^{\lambda}(X_{n})+\sqrt{2\gamma}\,\xi_{n+1}$ (8)
for $n\geq 0$, where $(\xi_{n})$ is a sequence of i.i.d. $d$-dimensional standard Gaussians.
@@
for $n\geq 0$, where $(\xi_{n})$ are i.i.d. standard Gaussians. The computational performance of the algorithm depends strongly on the capacity to evaluate efficiently $\operatorname{prox}_{U}^{\gamma}$.
@@
$\mathrm{d}\mathbf{X}_{\lambda,t}=-(\nabla g_{1}(\mathbf{X}_{\lambda,t})+\nabla g_{2}^{\lambda}(\mathbf{X}_{\lambda,t}))\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t},$
@@
$\mathrm{d}\mathbf{X}_{\lambda,t}=-\nabla g_{1}(\mathbf{X}_{\lambda,t})+\lambda^{-1}(\operatorname{prox}_{g_{2}}^{\lambda}(\mathbf{X}_{\lambda,t})-\mathbf{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}.$
@@ 2.3.2 Proximal gradient MCMC methods
where $(\xi_{n})_{n\in\mathbb{N}}$ is a sequence of i.i.d. $d$-dimensional standard Gaussians and $\gamma$ is the time discretisation step. Combining all the steps we obtain the proximal gradient Langevin algorithm (PGLA)
@@ 3 Proximal Interacting Particle Methods for MMLE
$\operatorname{prox}_{U}^{\lambda}(\theta,x)$ $=(\operatorname{prox}_{U}^{\lambda}(\theta,x)_{\theta},\operatorname{prox}_{U}^{\lambda}(\theta,x)_{x})$
$=\operatorname*{arg\,min}_{z_{0}\in\mathbb{R}^{d_{\theta}},z\in\mathbb{R}^{d_{x}}}\;\left\{U(z_{0},z)+\|(z_{0},z)-(\theta,x)\|^{2}/(2\lambda)\right\}.$
@@ 4.1 Assumptions
$\langle v-v^{\prime},\nabla U^{\lambda}(v)-\nabla U^{\lambda}(v^{\prime})\rangle$ $=\langle v-v^{\prime},\nabla g_{1}(v)-\nabla g_{1}(v^{\prime})\rangle+\langle v-v^{\prime},\nabla g_{2}^{\lambda}(v)-\nabla g_{2}^{\lambda}(v^{\prime})\rangle$
$\geq\mu\|v-v^{\prime}\|^{2}.$
@@ 4.2 The proof strategy
$\pi^{N}_{\lambda,\Theta}(\theta)\propto\int_{{}^{d_{x}}}...\int_{{}^{d_{x}}}e^{-\sum^{N}_{i=1}U^{\lambda}(\theta,x_{i})}\mathrm{d}x_{1}\mathrm{d}x_{2}\dots\mathrm{d}x_{N}=\biggr{(}\int_{{}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x\biggr{)}^{N}.$
@@
The first term is a concentration result, characterising the concentration of the $\theta$-marginal of the target measure of the SDE (9)–(10) around the maximiser of $p_{\theta}(y)$. Handling this term is not trivial since the maximisers of $p_{\theta}(y)$ and of $p_{\theta}^{\lambda}(y):=\int_{{}^{d_{x}}}p_{\theta}^{\lambda}(x,y)\mathrm{d}x$ are not necessarily the same as we clarify in the next section. The second term is a convergence term, which characterises the convergence of the solution of the SDE to its target measure. The last term is a discretisation term, which characterises the error introduced by discretising the SDE. By separately bounding each term, we provide nonasymptotic results for the convergence of the proximal interacting particle algorithms. Proofs of the following results are provided in Appendices A and B.
@@ 4.3 Nonasymptotic analysis of MYIPLA
$\mathbb{E}[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq$ $\frac{\lambda}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}+\sqrt{\frac{d_{\theta}}{N\mu}}$
$+e^{-\mu n\gamma}\bigg{(}\mathbb{E}[\|Z_{0}^{N}-z_{\star}\|^{2}]^{1/2}+\Big{(}\frac{d_{x}N+d_{\theta}}{N\mu}\Big{)}^{1/2}\bigg{)}$
@@
$\|\bar{\theta}_{\star}-\bar{\theta}_{\star,\lambda}\|\leq\frac{\lambda}{\mu}\left(\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\right)+\mathcal{O}(\lambda^{2}),$
with $A,B$ given in A5. This term quantifies the distance between maximisers of $p_{\theta}(y)$ and $p^{\lambda}_{\theta}(y)$. Next, consider the stationary measure of the SDE (9)–(10) denoted by $\pi_{\lambda,\star}^{N}(\theta,x_{1},\ldots,x_{N})$. We are interested in the $\theta$-marginal of this measure, denoted by $\pi_{\lambda,\Theta}^{N}$. Using a form of the Prékopa-Leindler inequality for strong convexity [Saumard and Wellner, 2014, Theorem 3.8], $\pi_{\lambda,\Theta}^{N}$ is $N\mu$-strongly log-concave and by Lemma A.8 of Altschuler and Chewi [2023]
$W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\star}})\leq\sqrt{\frac{d_{\theta}}{N\mu}},$
@@ 4.4 Nonasymptotic analysis of PIPGLA
To derive a convergence rate for PIPGLA we make the following additional assumption.
@@
We assume that $\|\nabla^{0}g_{2}(\theta,x)\|^{2}\leq C$ for all $\theta\in\mathbb{R}^{d_{\theta}}$ and $x\in\mathbb{R}^{d_{x}}$.
This assumption is similar to Salim et al. [2019, Assumption 5] without involving additional random variables.
@@
$\mathbb{E}\left[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}\right]^{1/2}\leq$ $\sqrt{\frac{d_{\theta}}{N\mu}}+\frac{\lambda^{n/2}(1-\gamma\mu)^{n/2}}{\gamma^{n/2}}W_{2}(\mathcal{L}({Z_{0}^{N}}),\pi^{N})$ (20)
$+\left(\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda NC)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}\right)^{1/2},$
@@
Similarly to the previous result, we can split the errors as follows:
$W_{2}(\mathcal{L}({\theta}^{N}_{n}),\delta_{\bar{\theta}_{\star}})\leq W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\Theta}^{N})+W_{2}(\pi_{\Theta}^{N},\mathcal{L}({\theta_{n}^{N}})).$
Using a similar result to Altschuler and Chewi [2023, Lemma 8], the concentration term $W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\Theta}^{N})$ is bounded by
$W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\Theta}^{N})\leq\sqrt{\frac{d_{\theta}}{N\mu}}.$
We derive novel nonasymptotic bounds for the other error term (see Corollary A.6 in Appendix A.4). Intuitively, $W_{2}(\pi_{\Theta}^{N},\mathcal{L}({\theta_{n}^{N}}))$ controls the convergence to $\pi_{\Theta}^{N}$ and the error due to time discretisation.